# Note on two-point mean square displacement

### v1

@@ III.1 Relative vector analysis
Comparing Eqs. (26) and (34), one find ${\rm MSCV}(\tau)=2{\rm MSD}(\tau)$ in short time-scale regime ($\tau\ll\tau_{m}$), the same relation to the case of independent probes (Sec. II.1), see Fig. 4 (e).
III.2 Distance analysis
We will follow the steps illustrated in Sec. II.2 for the case of independent probes. We first need to find $P_{{\vec{d}}}({\vec{d}},\tau|{\vec{d}}_{0})$; the probability density function of the relative vector ${\vec{d}}$ at $t=t_{1}+\tau$ given the condition $|{\vec{d}}(t=t_{1})|=d_{0}$ in stationary state. Let $t_{0}\rightarrow-\infty$ in Eq. (21), the system is thus in equilibrium at $t=t_{1}$. From this equilibrium ensemble, we pick up a sub-ensemble filtered by the condition ${\vec{d}}(t_{1})=d_{0}/\sqrt{3}(1,1,1)$. From here on, we measure the time as a function of $\tau=t-t_{1}$ since we are interested in the subsequent time evolution after the initial condition at $t=t_{1}$. The probability density function takes the form
@@ III.2 Distance analysis
$g(\tau)$ $=$ [$\langle\left[\{d_{\alpha}(\tau)-d_{\alpha}(0)\}+\{d_{\alpha}(0)-\langle d_{\alpha}(\tau)\rangle\}\right]^{2}\rangle$ (37)
@@
×$\times\left[\exp{\left(-\frac{(d-d_{0}{\tilde{h}}(\tau))^{2}}{2g(\tau)}\right)}-\exp{\left(-\frac{(d+d_{0}{\tilde{h}}(\tau))^{2}}{2g(\tau)}\right)}\right]$
Figure 4 (a) or (b) show the time evolution of $P_{d}(d,\tau|d_{0})$ starting from a shrunk ($d_{0}/\sqrt{ma^{2}}<1$) or swollen ($d_{0}/\sqrt{ma^{2}}>1$) conformation, respectively. Unlike for the case of independent probes, the distributions in both cases approach towards the equilibrium distribution given by Eq. (39) after $\tau\simeq\tau_{m}$. From this, we obtain the MSCD conditioned on $d(0)=d_{0}$, see Fig. 4 (c);
@@
Comparing Eqs. (40) and (41), respectively, with Eqs. (8) and (9), it is clear how the connectivity effect comes in the statistical time evolution of inter-probe distance.
Experimentally, as discussed in Sec. II.2, more easily accessible quantity is the time averaged MSCD, see Eq. (13), obtained by averaging over the equilibrium distribution of $d_{0}$ given by Eq. (39). After performing the averaging operation, we find
@@
Comparing Eq. (45) with Eq. (26), we find ${\rm MSCD}(\tau)=2/3{\rm MSD}(\tau)$ in short time-scale regime ($\tau\ll\tau_{m}$), the same relation to the case of independent probes (Sec. II.2), see Fig. 4 (e).
IV Discussions
IV.1 General property of MSCV and MSCD
@@
In Sec. II, we have seen the relation
${\rm MSD}(\tau)=\frac{1}{2}{\rm MSCV}(\tau)=\frac{3}{2}{\rm MSCD}(\tau)$ (46)
for two independently moving probes, see also Eqs. (18) and (19) for the analysis in the center-of-mass frame (Sec. II.3). The analysis of Rouse model in Sec. III.2 shows that the same relation also holds for the intramolecular pair of probes in short-time scale ($\tau\ll\tau_{m}$). Although our analyses in these preceding sections are based on simple models, i.e., Brownian motion model in Sec. II.2 and Rouse model in Sec. III.2, we now discuss that the relation (46) is quite generally expected even for more complex situations, i.e., probe particles performing anomalous diffusion (see, for instance, Fig. 2 (b)) and probes belonging to more general polymers, etc., as a consequence of defining properties of MSCV and MSCD.
First, the relation between MSD and MSCV can be derived from the following simple argument. Since $\Delta{\vec{d}}(\tau)=\Delta{\vec{r}}_{1}(\tau)-\Delta{\vec{r}}_{2}(\tau)$, one can transform Eq. (3) as
@@
For the relation between MSD and MSCD, we start from a relation ${\vec{d}}(t_{0}+\tau)={\vec{d}}(t_{0})+\Delta{\vec{r}}_{1}(\tau)-\Delta{\vec{r}}_{2}(\tau)$, and introduce a unit vector ${\hat{u}}$ associated with the relative vector at $t=t_{0}$, i.e., ${\vec{d}}(t_{0})=d_{0}{\hat{u}}$, see also Refs. Ponisch2018 ; Yesbolatova2022 for a related discussion. Then, the distance at time $t=t_{0}+\tau$ is expressed as
@@
This generalizes Eq. (12); in long-time regime ($\kappa\gg 1$), ${\rm MSCD}(\tau|d_{0})$ is the same as ${\rm MSCV}(\tau)$ and equal to twice ${\rm MSD}(\tau)$. On the other hand, a reduction factor $1/3$, or more generally, 1/(space dimension), appears in short-time regime ($\kappa\ll 1$), since we only count one component, i.e., along ${\vec{d}}(t_{0})$ in the calculation. As we have seen in Sec. II.2, this short-time scale behavior dominates after averaging over $d_{0}$ for a pair of independent probes in stationary state. Although the pair of intramolecular probes analyzed in Sec. III.2 belong to the same chain, their behaviors are essentially independent in the short-time scale $\tau\ll\tau_{m}$, as Eq. (22) implies. Hence, in both cases, one expect the relation (46). On the other hand, a pair of intramolecular probes is no longer independent of each other at $\tau\gtrsim\tau_{m}$, hence, ${\rm MSCD}(\tau)$ as well as ${\rm MSCV}(\tau)$ eventually approach a $m$-dependent constant value.
In Fig. 5, we show an example of MSCD analysis applied to the trajectory data obtained from molecular dynamic simulation of entangled polymer solutions. Here we employ a rather standard model of polymer solution with the chain length $N=600$ and the monomer volume fraction $\phi=0.1$ (see Appendix B for more details). As shown, MSD of tagged monomer exhibits rich anomalous behaviors, i.e., early $\tau^{1/2}$, subsequent slowing-down, a signature of the reptation dynamics (${\rm MSD}\sim\tau^{1/4}$) that is followed by later $\tau\sim t^{1/2}$ scaling DoiEdwards ; de_Gennes_scaling . In addition to MSD, we have calculated ${\rm MSCD}$ using probes on different polymers (see Fig. 2 (b)) and on the same polymer (see Fig. 2 (d)). We have also calculated ${\rm MSCD_{c}}$ from the distance analysis between a probe and the center of mass of the system (see Fig. 2 (c)). It is remarkable that all these quantities can be mapped to MSD quite accurately according to the formulas (46) and (19).
IV.2 Motional correlation
The above discussion clarifies that the relation (46) among MSCV, MSCD and MSD is generally expected for a pair of independent probes. In turn, this indicates that two-point MSD analysis can be combined with the standard MSD analysis to quantify the degree of correlation between motion of two probes. We illustrate the idea using the MSCV as a two-point MSD output. Assuming equal mobility for two probes, Eq (LABEL:MSCV_eq) leads to the expression of displacement correlation between two probes in the time-scale $\tau$ in terms of ${\rm MSD}(\tau)$ and ${\rm MSCV}(\tau)$;
$W(m,\tau)=ma^{2}\left[{\rm erf}\left(\sqrt{\frac{\tau_{m}}{4\tau}}\right)+\sqrt{\frac{4\tau}{\pi\tau_{m}}}\exp{\left(-\frac{\tau_{m}}{4\tau}\right)}\right]$
In Fig. (6) (a), we plot $W(m,\tau)$ as a function of $\tau$ for three different inter-probe separations. The correlation between the probes’ displacement is almost zero in the short time-scale in all cases, but starts to rise at the characteristic time-scale $\tau\simeq\tau_{m}$. Two probes thus do not feel each other in the time-scale $\tau<\tau_{m}$, but move together in longer time-scale, in which $W(m,\tau)\sim{\rm MSD}(\tau)$. One can also look at $W(m,\tau)$ as a function of $m$ for a given $\tau$, or in the same way as a function of $R_{m}$ that represents the characteristic spatial distance between two tagged monomers $m$-apart along the chain. With $W(0,\tau)={\rm MSD}(\tau)$, the normalized quantity $W(m,\tau)/{\rm MSD}(\tau)$ represents the degree of motional correlation in the time-scale $\tau$ for two tagged monomers with spatial distance $R_{m}=am^{1/2}$ apart. In Fig. (6) (b), we plot $W(m,\tau)/{\rm MSD}(\tau)$ as a function of $R_{m}$ for three different time-scales. One can identify the characteristic “domain”, whose size grows with time-scale as $\sim a\tau^{1/4}$.
IV.3 Non-Gaussian parameter
@@
$\langle\Delta d(\tau)^{4}\rangle$ $\simeq$ $\langle(\Delta x_{1}(\tau)-\Delta x_{2}(\tau))^{4}\rangle$
$=$ $\langle\Delta x_{1}(\tau)^{4}\rangle+6\langle\Delta x_{1}(\tau)^{2}\Delta x_{2}(\tau)^{2}\rangle+\langle\Delta x_{2}(\tau)^{4}\rangle$
@@
We applied the analysis to the entangled linear polymer solution, in which the entanglement effect is known to cause non-Gaussian dynamics Shen2022 . Employing the same data used in Fig. 5 (chain length $N=600$, monomer volume fraction $\phi=0.1$), we calculated the distance-based non-Gaussian parameter from (i) independent probe pair, (ii) intramolecular probe pair, (iii) a single probe and the center of mass. The results are shown in Fig. 6. Compared to a standard non-Gaussian parameter in single particle tracking analysis
@@
V Summary
@@
A growing need for two-point MSD has been seen in several recent experiments, in particular, those tracking chromatin loci in living cellular nucleus Marshall1997 ; Mine-Hattab2012 ; Ochiai2015 ; Arai2017 ; Khanna2019 ; Yesbolatova2022 ; Gabriele2022 ; Bruckner2023 ; Nozaki2023 ; Minami2024 . In this note, we have summarized the idea of two-point MSD, the definition of two variants ${\rm MSCV}$ and ${\rm MSCD}$ and their basic properties for a pair of independent probes as well as intramolecular probes. As is clear from their definitions, ${\rm MSCD}$ is more suitable to quantify the intrinsic mobility of the probe when the rotational motion of the container comes in. Otherwise, ${\rm MSCV}$ can be used alike.
@@
Simple reasoning in Sect. IV.1 shows that the relation (46) among ${\rm MSCV}$, ${\rm MSCD}$ and ${\rm MSD}$ generally holds for a pair of independent probes moving in free space in the stationary state. For pairs that are not independent, on the other hand, the relation (46) breaks down. In that case, one can exploit this fact by combining two-point MSD measurement with ${\rm MSD}$, which enables us to extract and quantify the degree of motional correlation between those probe pairs (see Sec. IV.2).
@@ Acknowledgements
This work was supported by JSPS KAKENHI (Grant Nos. JP23H00369, JP23H04290 and JP24K00602). We thanks A. Kimura for discussion and motivating us to summarize this note.

### v2

@@ IV.1 Relative vector analysis
Comparing Eqs. (29) and (37), one find ${\rm MSCV}(\tau)=2{\rm MSD}(\tau)$ in short time-scale regime ($\tau\ll\tau_{m}$), the same relation to the case of independent probes (Sec. III.1), see Fig. 5 (e).
IV.2 Distance analysis
We will follow the steps illustrated in Sec. III.2 for the case of independent probes. We first need to find $P_{{\vec{d}}}({\vec{d}},\tau|{\vec{d}}_{0})$; the probability density function of the relative vector $\vec{d}$} at $t=t_{1}+\tau$ given the condition $|{\vec{d}}(t=t_{1})|=d_{0}$ in stationary state. Let $t_{0}\rightarrow$-$\infty$ in Eq. (23), the system is thus in equilibrium at $t=t_{1}$. From this equilibrium ensemble, we pick up a sub-ensemble filtered by the condition ${\vec{d}}(t_{1})=d_{0}/\sqrt{3}(1,1,1)$. From here on, we measure the time as a function of $\tau=t-t_{1}$ since we are interested in the subsequent time evolution after the initial condition at $t=t_{1}$. The probability density function takes the form
@@ IV.2 Distance analysis
$g(\tau)$ $=$ $\langle\left[\{d_{\alpha}(\tau)-d_{\alpha}(0)\}+\{d_{\alpha}(0)-\langle d_{\alpha}(\tau)\rangle\}\right]^{2}\rangle$ (40)
@@
$\times\left[\exp{\left(-\frac{(d-d_{0}{\tilde{h}}(\tau))^{2}}{2g(\tau)}\right)}-\exp{\left(-\frac{(d+d_{0}{\tilde{h}}(\tau))^{2}}{2g(\tau)}\right)}\right]$
Figure 5 (a) or (b) shows the time evolution of $P_{d}(d,\tau|d_{0})$ starting from a shrunk ($d_{0}/\sqrt{ma^{2}}<1$) or swollen ($d_{0}/\sqrt{ma^{2}}>1$) conformation, respectively. Unlike for the case of independent probes, the distributions in both cases approach towards the equilibrium distribution given by Eq. (42) after $\tau\simeq\tau_{m}$. From this, we obtain the MSCD conditioned on $d(0)=d_{0}$, see Fig. 5 (c);
@@
Comparing Eqs. (43) and (44), respectively, with Eqs. (LABEL:P_d_indep) and (11), it is clear how the connectivity effect comes in the statistical time evolution of inter-probe distance.
Experimentally, as discussed in Sec. III.2, more easily accessible quantity is the time averaged MSCD, see Eq. (15), obtained by averaging over the equilibrium distribution of $d_{0}$ given by Eq. (42). After performing the averaging operation, we find
@@
Comparing Eq. (48) with Eq. (29), we find ${\rm MSCD}(\tau)=2/3{\rm MSD}(\tau)$ in short time-scale regime ($\tau\ll\tau_{m}$), the same relation to the case of independent probes (Sec. III.2), see Fig. 5 (e).
V Discussions
V.1 General property of MSCV and MSCD
@@
In Sec. II.2, we have claimed the formula (5) for the relation between MSD, MSCV and MSCD. So far, we have verified it for two independently moving probes in Sec. III, see also Eqs. (20) and (21) for the analysis in the center-of-mass frame (Sec. III.3), and for the intramolecular pair of probes within a polymer in short-time scale ($\tau\ll\tau_{m}$) in Sec. IV.2. Although our analyses in these preceding sections are based on simple models, i.e., Brownian motion model in Sec. III and Rouse model in Sec. IV, we now discuss that the relation (5) is quite generally expected even for more complex situations, i.e., probe particles performing anomalous diffusion (see, for instance, Fig. 2 (b)) and probes belonging to more general polymers, etc., as a consequence of defining properties of MSCV and MSCD. For the latter, the discussion allows us to provide the precise definition of the short time scale mentioned in Sec. II.2, and clarify the underlying physics behind the formula (5).
We begin by repeating a simple argument in Sec. II.2 for the relation between MSD and MSCV. Since $\Delta{\vec{d}}(\tau)=\Delta{\vec{r}}_{1}(\tau)-\Delta{\vec{r}}_{2}(\tau)$, one can transform Eq. (3) as
@@
For the relation between MSD and MSCD, we start from a relation ${\vec{d}}(t_{0}+\tau)={\vec{d}}(t_{0})+\Delta{\vec{r}}_{1}(\tau)-\Delta{\vec{r}}_{2}(\tau)$, and introduce a unit vector $^{\hat{u}}$ associated with the relative vector at $t=t_{0}$, i.e., $^{\vec{d}}(t_{0})=d_{0}{\hat{u}}$, see also Refs. Ponisch2018 ; Yesbolatova2022 for a related discussion. Then, the distance at time $t=t_{0}+\tau$ is expressed as
@@
This generalizes Eq. (14); in long-time regime ($\kappa\gg 1$), ${\rm MSCD}(\tau|d_{0})$ is the same as ${\rm MSCV}(\tau)$ and equal to twice ${\rm MSD}(\tau)$. On the other hand, a reduction factor $1/3$, or more generally, 1/(space dimension), appears in short-time regime ($\kappa\ll 1$), since we only count one component, i.e., along ${\vec{d}}(t_{0})$ in the calculation. As we have seen in Sec. III.2, this short-time scale behavior dominates after averaging over $d_{0}$ for a pair of independent probes in stationary state. Although the pair of intramolecular probes analyzed in Sec. IV.2 belong to the same chain, their behaviors are essentially independent in the short-time scale $\tau\ll\tau_{m}$, as Eq. (24) implies. Hence, in both cases, one expect the relation (5). On the other hand, a pair of intramolecular probes is no longer independent of each other at $\tau>\tau_{m}$, hence, ${\rm MSCD}(\tau)$ as well as ${\rm MSCV}(\tau)$ eventually approach a $m$-dependent constant value.
In Fig. 6, we show an example of MSCD analysis applied to the trajectory data obtained from molecular dynamic simulation of entangled polymer solutions. Here we employ a rather standard model of polymer solution with the chain length $N=600$ and the monomer volume fraction $\phi=0.1$ (see Appendix B for more details). As shown, MSD of tagged monomer exhibits rich anomalous behaviors, i.e., early $\tau^{1/2}$, subsequent slowing-down, a signature of the reptation dynamics (${\rm MSD}\sim\tau^{1/4}$) that is followed by later ${\rm MSD}\sim\tau^{1/2}$ scaling DoiEdwards ; de_Gennes_scaling . In addition to MSD, we have calculated $\rm MSCD$} using probes on different polymers (see Fig. 2 (b)) and on the same polymer (see Fig. 2 (d)). We have also calculated $\rm MSCD_{c}$} from the distance analysis between a probe and the center of mass of the system (see Fig. 2 (c)). It is remarkable that all these quantities can be mapped to MSD quite accurately according to the formulas (5) and (21).
V.2 Non-Gaussian parameter
@@
$\langle\Delta d(\tau)^{4}\rangle$ $\simeq$ $\langle(\Delta x_{1}(\tau)-\Delta x_{2}(\tau))^{4}\rangle$ (58)
$=$ $\langle\Delta x_{1}(\tau)^{4}\rangle+6\langle\Delta x_{1}(\tau)^{2}\rangle\langle\Delta x_{2}(\tau)^{2}\rangle$
$+\langle\Delta x_{2}(\tau)^{4}\rangle$
@@
We applied the analysis to the entangled linear polymer solution, in which the entanglement effect is known to cause non-Gaussian dynamics Shen2022 . Employing the same data used in Fig. 6 (chain length $N=600$, monomer volume fraction $\phi=0.1$), we calculated the distance-based non-Gaussian parameter from (i) independent probe pair, (ii) intramolecular probe pair, (iii) a single probe and the center of mass. The results are shown in Fig. 7. Compared to a standard non-Gaussian parameter in single particle tracking analysis  [^1] This definition comes from the following observation. If the probe displacement is isotropic in space and follows Gaussian statistics, the following calculation holds; $\langle|\Delta{\vec{r}}(\tau)|^{4}\rangle=\langle[\Delta x(\tau)^{2}+\Delta y(\tau)^{2}+\Delta z(\tau)^{2}]^{2}\rangle=3\langle\Delta x(\tau)^{4}\rangle+6\langle\Delta x(\tau)^{2}\rangle^{2}=15\langle\Delta x(\tau)^{2}\rangle^{2}=5[{\rm MSD}(\tau)]^{2}/3$.
@@
V.3 Motional correlation
In the discussion of two-point MSD, one of the central quantities is the correlation in the motion of probes. Here, it is important to notice that there are two types of correlations, i.e., either extrinsic arising from substrate motion or intrinsic correlation between probes. In this paper, we have been mostly concerned with the correlation of extrinsic origin. Indeed, a basic idea of two-point MSD analysis is filtering out such a correlation arising from the substrate motion. Specifically, MSCD has the property ${\rm MSCD}(\tau)={\rm MSCD}^{(o)}(\tau)$, thus, can be calculated from $\Delta{\vec{r}}_{i}^{(o)}(\tau)$ for $i=1,2$ (see Fig. 3), with which one can obtain MSD using Eq. (5) as long as there is no intrinsic correlation in their dynamics. As a different, but related methodology, Oliveira et al. has recently proposed a framework that, through the analysis of correlated motion of probes, provides a way to estimate the substrate motion, giving an access to the intrinsic probe mobility Oliveira2021 .
In the remainder of this subsection, we will focus on the opposite situation, where the substrate motion is negligible, i.e., $\Delta{\vec{r}}^{(o)}(\tau)=\Delta{\vec{r}}(\tau)$, thus, not only MSCD but also MSD and MSCV can be calculated from $\Delta{\vec{r}}^{(o)}(\tau)$, but there is intrinsic correlation in the dynamics of probe pairs, invalidating the relation (5). In turn, this indicates that two-point MSD analysis can be combined with the standard MSD analysis to quantify the degree of intrinsic correlation between motion of two probes.
We illustrate the idea using the MSCV as a two-point MSD output. Assuming equal mobility for two probes, Eq (LABEL:MSCV_eq) leads to the expression of displacement correlation between two probes in the time-scale $\tau$ in terms of ${\rm MSD}(\tau)$ and ${\rm MSCV}(\tau)$;
$W(m,\tau)=ma^{2}\left[\sqrt{\frac{4\tau}{\pi\tau_{m}}}\exp{\left(-\frac{\tau_{m}}{4\tau}\right)}-{\rm erfc}\left(\sqrt{\frac{\tau_{m}}{4\tau}}\right)\right]$
where ${\rm erfc}(z)=1-{\rm erf}(z)$ is complementary error function. In Fig. (8) (a), we plot $W(m,\tau)$ as a function of $\tau$ for three different inter-probe separations. The correlation between the probes’ displacement is almost zero in the short time-scale in all cases, but starts to rise at the characteristic time-scale $\tau\simeq\tau_{m}$. Two probes thus do not feel each other in the time-scale $\tau<\tau_{m}$, but move together in longer time-scale, in which $W(m,\tau)\sim{\rm MSD}(\tau)$. One can also look at $W(m,\tau)$ as a function of $m$ for a given $\tau$, or in the same way as a function of $R_{m}$ that represents the characteristic spatial distance between two tagged monomers $m$-apart along the chain. With $W(0,\tau)={\rm MSD}(\tau)$, the normalized quantity $W(m,\tau)/{\rm MSD}(\tau)$ represents the degree of motional correlation in the time-scale $\tau$ for two tagged monomers with spatial distance $R_{m}=am^{1/2}$ apart. In Fig. (8) (b), we plot $W(m,\tau)/{\rm MSD}(\tau)$ as a function of $R_{m}$ for three different time-scales. One can identify the characteristic “domain”, whose size grows with time-scale as $\sim a\tau^{1/4}$.
VI Summary
@@
The growing need for two-point MSD has been seen in several recent experiments, in particular those that track chromatin loci in living cell nucleus Marshall1997 ; Mine-Hattab2012 ; Ochiai2015 ; Arai2017 ; Khanna2019 ; Yesbolatova2022 ; Gabriele2022 ; Bruckner2023 ; Nozaki2023 ; Minami2024 . Yet, the term two-point MSD has so far been used depending on context with nuances, lacking its precise definition. In this note, we have summarized the idea of two-point MSD, the definition of two variants $\rm MSCV$} and $\rm MSCD$} and their basic properties for a pair of independent probes as well as intramolecular probes. As is clear from their definitions, $\rm MSCD$} is more suitable to quantify the intrinsic mobility of the probe when the rotational motion of the substrate comes in. Otherwise, $\rm MSCV$} can be used alike, see the demonstration in Sec. II.1. We note that a apparently similar set-up has been conventionally employed in micro-rheology experiments but with a different motivation. Indeed, while the two-point micro-rheology is mainly motivated to eliminate the influence of local inhomogeneities in complex media Crocker2000 , the two-point MSD discussed here aims at providing a way to quantify the intrinsic mobility of probes by eliminating the influence of the extrinsic motion.
@@
Simple reasoning in Sect. V.1 shows that the relation (5) among $\rm MSCV$}, $\rm MSCD$} and $\rm MSD$} generally holds for a pair of independent probes moving in free space in stationary state (or in systems with stationary increment). For pairs that are not independent, on the other hand, the relation (5) breaks down. In that case, one can exploit this fact by combining two-point MSD measurement with $\rm MSD$}, which enables us to extract and quantify the degree of motional correlation between those probe pairs (see Sec. V.3).
@@ Acknowledgements
This work was supported by JSPS KAKENHI (Grant Nos. JP23H00369, JP23H04290 and JP24K00602) and AGU Future Eagle Project. We thank A. Kimura for discussion and motivating us to summarize this note, and C. Lim for insightful comments on our note.